OpenStudy (zzr0ck3r):

E = {x/2^y:x,y elemnts of N} prove inf(E) = 0. (please use contrapositive)

OpenStudy (kinggeorge):

So I'm going to assume that \(0\in\mathbb{N}\), otherwise I don't think this could be true.

OpenStudy (kinggeorge):

I was wrong, we don't have to assume 0 is in the natural numbers. There is a more elegant way. So we need to start by assuming \(\inf(E)\neq0\). We have two cases Case 1: Suppose \(\inf(E)<0\). That means that there exists some \(a\in E\) such that \(a<0\). However, \(x\) and \(2^y\) are always non-negative since \(\mathbb{N}\) is non-negative. This is a contradiction, so \(\inf(E)\geq0\). Case 2: Let \(a=\inf(E)>0\). Then \(a>0\), and given any \(b\in E\), \(a\le b\). Now, let \[\Large a=\frac{r}{2^s}\]Notice that \[\Large \frac{r}{2^{s+1}}< \frac{r}{2^s}\]Hence, \(a\) is not \(\inf(E)\), so \(\inf(E)\leq 0\). It follows that \(\inf(E)=0\).

OpenStudy (kinggeorge):

Let me know if you need help understanding certain steps.

OpenStudy (zzr0ck3r):

word, good idea. TY